The perimeter of a rectangle is 168 inches. The length exceeds the width by 62 inches. Find the length and width.
Answer by jonvaliente(64)

Let w = width of the rectangle
Let w+62 = length of the rectangle
P = 2L + 2W
168 = 2*(w+62) + 2*(w)
168 = 2w + 124 + 2w
168-124 = 4w + 124 - 124
44 = 4w
w = 11
length = w+62=11+62=73
Therefore, the width is 11 and the length is 73.